# A parameterized generative adversarial network using cyclic projection for explainable medical image classifications

### 07 Dec 2023 v2

@@ 1 Introduction
Deep neural networks have achieved success in computer vision fields [1]. In these fields, a large-scale dataset is crucial for effectively training of neural networks. However, neural networks trained on small-scale datasets can lead to overfitting and poor generalization. Although several regularization techniques [2, 3, 4, 5] have been developed to prevent overfitting, augmenting training data is an effective option to address the data insufficiency.
Researchers have developed two main data augmentation methods in the past decade. Conventional augmentation method can explore prior knowledge [6, 7], including random cropping, rotations, flipping and erasing, and adding noise. However, it depends on domain knowledge and the intra-domain augmented data contributes little to the description of decision boundaries for downstream classifications. Generative adversarial network (GAN) [8] and its variations have also emerged as augmentation means [9, 10, 11]. However, the synthetic images are quality uncertain and variety insufficiency because of the mode collapsing [12]. Moreover, there is still uncertainty regarding the domain labels of the synthetic images even if soft labels are used [13, 14].
@@
To overcome these issues, we propose a parameterized GAN (ParaGAN) enabling controlling the degree variation of synthetic images. This is achieved by leveraging target-domain samples’ distances to the optimal hyperplane as controllable parameters in forward path and considering source images’ distances to the boundary in the backward reconstruction path. Experiments show that ParaGAN augmented method consistently outperforms the downstream classifier and the state of the arts on limited datasets, and provides a more transparent explanation than using Grad-CAM [21].
@@ 2.1 Dataset Acquisition and Evaluation Metrics
We conduct experiments over multiple public datasets: Breast Ultrasound Images Dataset (BUSI) [22], Breast Ultrasound Images from UDIAT Diagnostic Center (UDIAT)[23] and CT Scan Dataset about COVID-19 (COVID-CT) [24]. The BUSI and UDIAT datasets are mixed. We perform evaluations with two metrics: accuracy (ACC) and the area under the receiver operating characteristic curve (AUC).
@@ 2.2 Parameterized Generative Adversarial Network
$\mathcal{L}_{\rm hinge}(t,\hat{t})=\frac{1}{N}\sum_{i=1}^{N}max[0,1-t_{i}(w^{T}z_{i}+b))],$ (1)
We develop parameterized generators $G_{x2y}$ and $G_{y2x}$ taking the projection distances to hyperplane as input parameters. To achieve this, we use pre-trained $C_{aux}$ to measure projection distances from the real images to the optimal hyperplane ($w^{T}z+b=0$). Given two random real images $x$ and $y$ in domain $X$ and domain $Y$ respectively, their vertical projection distances $d_{x}$ and $d_{y}$ are as follows:
$d_{x}=C_{aux}(x),x\in X;d_{y}=C_{aux}(y),y\in Y.$ (2)
Figure 1(b) illustrates the training process of ParaGAN. Firstly, $G_{x2y}$ translates a source image $x$ into $G_{x2y}(x,+d_{y})$ and $D_{Y}$ distinguishes between this translated image and the real image $y$, and vice versa for $G_{y2x}$ and $D_{X}$. Secondly, an auxiliary classifier $C_{aux}$ reconstructs the vertical projection distances $d_{y}$ and $d_{x}$ from the translated images $G_{x2y}(x,+d_{y})$ and $G_{y2x}(y,-d_{x})$, respectively. Thirdly, the source image $x$ is reconstructed from the translated image $G_{x2y}(x,+d_{y})$ using its projection distance $d_{x}$, and vice versa for source $y$.
@@ 2.3 Objective Function
$\begin{split}\mathcal{L}_{\rm GAN}(G_{x2y},&D_{Y},X,Y)=\mathbb{E}_{y\in Y}[logD_{Y}(y)]\\ &+\mathbb{E}_{x\in X}[log(1-D_{Y}(G_{x2y}(x,+d_{y}))],\end{split}$ (3)
@@
$\begin{split}\mathcal{L}_{\rm proj}(&G_{x2y},G_{y2x},C_{aux})\\ &=\mathbb{E}_{x\in X}[||(+d_{y})-C_{aux}(G_{x2y}(x,+d_{y}))||_{2}^{2}]\\ &+\mathbb{E}_{y\in Y}[||(-d_{x})-C_{aux}(G_{y2x}(y,-d_{x}))||_{2}^{2}].\end{split}$ (4)
Cycle consistency is introduced to establish relationships between individual input $x_{i}$ and a desired output $y_{i}$. Thus, the cycle consistency loss is formulated as:
$\begin{split}\mathcal{L}_{\rm cyc}(&G_{x2y},G_{y2x})\\ &=\mathbb{E}_{x\in X}[||x-G_{y2x}(G_{x2y}(x,+d_{y}),-d_{x})||_{1}]\\ &+\mathbb{E}_{y\in Y}[||y-G_{x2y}(G_{y2x}(y,-d_{x}),+d_{y})||_{1}].\end{split}$ (5)
@@
$\begin{split}\mathcal{L}_{\rm ParaGAN}(&G_{x2y},G_{y2x},C_{aux},D_{X},D_{Y})\\ &=\mathcal{L}_{\rm GAN}(G_{x2y},D_{Y},X,Y)\\ &+\mathcal{L}_{\rm GAN}(G_{y2x},D_{X},Y,X)\\ &+\lambda_{\rm proj}\mathcal{L}_{\rm proj}(G_{x2y},G_{y2x},C_{aux})\\ &+\lambda_{\rm cyc}\mathcal{L}_{\rm cyc}(G_{x2y},G_{y2x}).\end{split}$ (6)
2.4 Downstream Classifier Optimization
$\mathcal{L}_{\rm C}(c,\hat{c})=\mathcal{L}_{\rm hinge}(c,\hat{c}_{real})+\alpha\mathcal{L}_{\rm hinge}(c,\hat{c}_{syn}).$ (7)
@@ 3.2 Training Settings
We train all networks with a learning rate of 0.0002 for the first 25 epochs and linearly decay the learning rate to 0 over the next 25 epochs. The hyperparameter $\lambda_{\rm cyc}$ is set to 10, and $\lambda_{\rm proj}$ is set to 0.1 considering the large scale of the distances. We adopt transfer learning for pretraining auxiliary classifier and training downstream classifiers using ImageNet Dataset. The auxiliary classifier $C_{aux}$ is not updated during the ParaGAN training procedure. We report ACC and AUC averaged over three runs.

### 14 Dec 2023 v3

@@ 1 Introduction
Deep neural networks have achieved success in computer vision fields [1], where a large-scale dataset is crucial for effectively training. However, training neural networks on the small-scale datasets leads to the overfitting and poor generalization. Although regularization techniques [2, 3, 4, 5] have been developed to prevent overfitting, augmenting training data is an effective option to address the data insufficiency.
Researchers have developed two main data augmentation methods in the past decade. Conventional augmentation method can explore prior knowledge [6, 7], including random cropping, flipping, etc. However, its intra-domain augmented data contributes little to the description of decision boundaries for downstream classifications. Generative adversarial network (GAN) [8] and its variations have also emerged as augmentation means [9, 10, 11]. However, the synthetic images are quality uncertain and variety insufficiency because of the mode collapsing [12]. Moreover, there is still uncertainty regarding the domain labels of the synthetic images even if soft labels are used [13, 14].
@@
To overcome these issues, we propose a parameterized GAN (ParaGAN) enabling controlling the degree variation of synthetic images. This is achieved by leveraging target-domain samples’ distances to the optimal hyperplane as controllable parameters in forward path and considering source images’ distances to the hyperplane in the backward reconstruction path. Experiments show that ParaGAN consistently outperforms the state-of-the-arts and provides a more transparent explanation than Grad-CAM [21].
@@ 2.1 Dataset Acquisition and Evaluation Metrics
We collect a mixed breast ultrasound datasets which contains BUSI [22] and UDIAT [23] and a COVID-19 Dataset (COVID-CT) [24]. Accuracy (ACC) and the area under the receiver operating characteristic curve are used for evaluation.
@@ 2.2 Parameterized Generative Adversarial Network
$\small\mathcal{L}_{\rm hinge}(t,\hat{t})=\frac{1}{N}\sum_{i=1}^{N}max[0,1-t_{i}(w^{T}z_{i}+b))],$ (1)
We develop parameterized generators $G_{X2Y}$ and $G_{Y2X}$ taking the projection distances to hyperplane as input parameters. To achieve this, we use pre-trained $C_{aux}$ to measure projection distances from the real images to the optimal hyperplane ($w_{aux}^{T}z+b_{aux}=0$). Given two random real images $x$ and $y$, their vertical projection distances $d_{x}$ and $d_{y}$ are as:
$\small d_{x}=\lvert C_{aux}(x)\rvert,x\in X;d_{y}=\lvert C_{aux}(y)\rvert,y\in Y.$ (2)
Figure 1(b) illustrates the synthesis phase. Firstly, $G_{X2Y}$ translates a source image $x$ into $G_{X2Y}(x,+d_{y})$ and $D_{Y}$ distinguishes between this translated image and the real image $y$, and vice versa for $G_{Y2X}$ and $D_{X}$. Secondly, an auxiliary classifier $C_{aux}$ reconstructs the vertical projection distances $d_{y}$ and $d_{x}$ from the translated images $G_{X2Y}(x,+d_{y})$ and $G_{Y2X}(y,-d_{x})$, respectively. Thirdly, the source image $x$ is reconstructed from the translated image $G_{X2Y}(x,+d_{y})$ using its projection distance $d_{x}$, and vice versa for source $y$.
@@ 2.3 Objective Function
$\small\begin{split}\mathcal{L}_{\rm GAN}(G_{X2Y},&D_{Y},X,Y)=\mathbb{E}_{y\in Y}[logD_{Y}(y)]\\ &+\mathbb{E}_{x\in X}[log(1-D_{Y}(G_{X2Y}(x,+d_{y}))],\end{split}$ (3)
@@
$\small\begin{split}\mathcal{L}_{\rm proj}(&G_{X2Y},G_{Y2X},C_{aux})\\ &=\mathbb{E}_{x\in X}[||(+d_{y})-C_{aux}(G_{X2Y}(x,+d_{y}))||_{2}^{2}]\\ &+\mathbb{E}_{y\in Y}[||(-d_{x})-C_{aux}(G_{Y2X}(y,-d_{x}))||_{2}^{2}].\end{split}$ (4)
Cycle consistency loss is introduced to establish relationships between individual input $x_{i}$ and a desired output $y_{i}$:
$\small\begin{split}\mathcal{L}_{\rm cyc}(&G_{X2Y},G_{Y2X})\\ &=\mathbb{E}_{x\in X}[||x-G_{Y2X}(G_{X2Y}(x,+d_{y}),-d_{x})||_{1}]\\ &+\mathbb{E}_{y\in Y}[||y-G_{X2Y}(G_{Y2X}(y,-d_{x}),+d_{y})||_{1}].\end{split}$ (5)
$\small\begin{split}&\mathcal{L}_{\rm ParaGAN}(G_{X2Y},G_{Y2X},C_{aux},D_{X},D_{Y})\\ &=\mathcal{L}_{\rm GAN}(G_{X2Y},D_{Y},X,Y)+\mathcal{L}_{\rm GAN}(G_{Y2X},D_{X},Y,X)\\ &+\lambda_{\rm proj}\mathcal{L}_{\rm proj}(G_{X2Y},G_{Y2X},C_{aux})\\ &+\lambda_{\rm cyc}\mathcal{L}_{\rm cyc}(G_{X2Y},G_{Y2X}).\end{split}$ (6)
2.4 Downstream Classifier Optimization
$\small\mathcal{L}_{\rm C}(c,\hat{c})=\mathcal{L}_{\rm hinge}(c,\hat{c}_{real})+\alpha\mathcal{L}_{\rm hinge}(c,\hat{c}_{syn}).$ (7)
@@ 3.2 Training Settings
We train all networks with a learning rate of 0.0002 for the first 25 epochs and linearly decay the learning rate to 0 over the next 25 epochs. The hyperparameter $\lambda_{\rm cyc}$ is set to 10, and $\lambda_{\rm proj}$ is set to 0.1 considering the large scale of the distances. We adopt transfer learning for training all classifiers using ImageNet Dataset. The auxiliary classifier $C_{aux}$ is not updated during the ParaGAN training procedure. We report ACC and AUC averaged over three runs.
@@ 5 Conclusion and Future Works
Acknowledgments
This work is supported in part by the Macao Polytechnic University Grant (RP/FCA-05/2022) and in part by the Science and Technology Development Fund of Macao (0041/2023/RIB2).